OpenStudy (ttop0816):

easy question! If a man earns $420 interest in one year at 7%, what was the amount of money invested? $29.40 $60.00 $600.00 $6,000.00

OpenStudy (agent0smith):

If x is the amount invested...\[\Large x \times \frac{ 7 }{ 100 } = 420\]can you solve for x?

OpenStudy (ttop0816):

0.6?

OpenStudy (agent0smith):

No... but show your work of how you got 0.6

OpenStudy (accidentalaichan):

\[420 \div .7\]

OpenStudy (jack1):

\[\frac{ 7\times x }{100 } = 420 \rightarrow x=\frac{ 420\times100 }{ 7 }\]

OpenStudy (agent0smith):

@ttop0816 show us your work so we can see where you went wrong.

OpenStudy (nbester98):

420 x .07 = 60. There's the answer

OpenStudy (agent0smith):

@nbester98 that's not correct. Read the context of the question... 7% of $60 is not $420.

OpenStudy (jack1):

lol @nbester98 , show me that bank ;D put in $60, get a 700% return within a year

OpenStudy (nbester98):

Typo and honest mistake plus too busy with other things. To get the accurate answer, multiply each of the four possible answers by .07 which should answer the question. That's what I was meant to say!

OpenStudy (nbester98):

@agent0smith @Jack1 trust my above answer satisfies.

OpenStudy (nbester98):

Typo and honest mistake plus too busy with other things. To get the accurate answer, multiply each of the four possible answers by .07 which should answer the question. That's what I was meant to say!

OpenStudy (agent0smith):

"multiply each of the four possible answers by .07 which should answer the question." That works here since it's multiple choice, but it's still good to know how to answer the question without multiple choice answers, in which case he'd need to use 0.07x = 420 to solve it.

OpenStudy (jack1):

@nbester98 all good dude, just light-hearted dig, not tryin to be mean hey and yeah, you're right in saying "test the answers by multiplying by .07" to see which comes to 420 is one way to solve it